OpenStudy (kainui):

When you have a distance it's a meter. An area is a meter squared. A volume is a meter cubed. So when you have equations with position, velocity, and acceleration, what do the higher powers of time represent? I know what a second is, but what is a second squared mean on its own when it's not m/s^2 ?

OpenStudy (anonymous):

Oh lol! I had the exact same question a while ago, look! http://openstudy.com/study#/updates/4fcc4910e4b0c6963ad7634b, any how, in my own words with acceleration and such meters per second squared is the time of the thing accelerating to THAT time.

OpenStudy (anonymous):

Oh, and staying to that question a second square on its own would be just an exponent, exponential time only applies to M/s2.

OpenStudy (anonymous):

delete the apostrophe.

OpenStudy (anonymous):

Hope that helped, IM not completely shure although, so dont trust me!

OpenStudy (kainui):

Sure, I know that acceleration is a change in velocity, but I'm not sure I understand your answer, explain it differently or something, it kind of makes sense.

OpenStudy (anonymous):

Ok!

OpenStudy (kainui):

I guess I'm interested in understanding what a second squared is when it is in the numerator and alone. Just a pure, second squared.

OpenStudy (anonymous):

Well

OpenStudy (anonymous):

From my knowledge of science a all, in the matter world I guess a second squared would be time duplicating itself.

OpenStudy (kainui):

So any time something accelerates, time duplicates? Haha.

OpenStudy (anonymous):

I guess you must have misunderstood, thats the only thing I can think of other THEN acceleration.

OpenStudy (kainui):

Ahh ok. Maybe I'm trying to make sense of something that really has no meaning.

OpenStudy (kainui):

Well, that can't be right though...

OpenStudy (anonymous):

I really doubt it, Like I said, dont trust me!

OpenStudy (kainui):

I'm looking back at understanding concepts like hertz and wavenumbers, just their simple units meanings since they represent inverse meters and seconds. It's just confusing. I guess I'm curious as to why units are so confusing to comprehend!

OpenStudy (anonymous):

e.g. acceleration is m/s^2. It's non-intuitive, but it helps to think about the definition of acceleration as the instantaneous rate of change in velocity (m/s) over time. I recommend lecture 2, mit's 8.01 physics... around 26:00.

OpenStudy (kainui):

I know what acceleration is, that really isn't the issue here. I am just trying to understand if a second squared means anything on its own or not. It's kind of a more complicated question.

OpenStudy (anonymous):

There is no physical significance to a square second.

OpenStudy (anonymous):

There are many terms in formulas that themselves don't mean anything. e.g. Force = kg*m/s^2, what's a kilogram-meter? Or to get square seconds in the numerator, how about \[mass = \frac{force}{acceleration}\] \[[1kg = 1\frac{(N)(s^2)}{m} \rightarrow 1s^2 = 1\frac{(kg)(m)}{N^2}\] So there's a definition of what a square second is, but I don't think it clears things up any. :-D